Consider the following two stocks: Stock A Stock B Expected Rtrn Variance 1196 6% 0.0324 0.0196 If the cortelation of their return is equal to 0.3, what is the standard deviation of a portfolico of the two stocks whose weight in stock A is equal to 0.3? (Answers rounded to two decimal digits)

hi guys how can i solve this
Answer #1

Weight of Stock A = W(A) = 0.3

Weight of Stock B = W(B) = 0.7

ER(A) = 11%

ER(B) = 6%

Var (A) = 0.0324

Var (B) = 0.0196

Correlation = 0.3

Covariance = Correlation x Var (A) x Var (B)

Cov = 0.3 x 0.0324 x 0.0196 = 0.00019051

Portfolio Variance = W(A)2 x Var(A) + W(B)2 x Var (B) + 2xW(A)xW(B)xCov(A,B)

Portfolio Variance = (0.3)2 x 0.0324 + (0.7)2 x 0.0196 + 2x0.3x0.7x0.00019051

Portfolio Variance = 0.002916 + 0.009604 + 0.00008001

Portfolio Variance = 0.0126

Standard deviation of portfolio = Square root of portfolio variance

Standard deviation of Portfolio = Sq. root (0.0126) = 0.11225